OpenStudy (anonymous):

John and Sue, ages eight and ten respectively, own a puppy named Jake. It is hared to take Jake for a walk that they devise a scheme to use two leashes. If John uses 24lbs of force and is at a -15 and Sue uses 27lbs of force at 18degrees. How hard is Jake pulling if the puppy is holding the children at a standstill? In other words find the resultant!!!

OpenStudy (anonymous):

poor doggy, being dragged :((

OpenStudy (anonymous):

is it difficult?

OpenStudy (anonymous):

i think its with vectors

OpenStudy (anonymous):

Do you know how to change vectors into their component forms?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Change the vectors into their component forms. Then you can add them.

OpenStudy (anonymous):

what are the vectors?

OpenStudy (anonymous):

"24lbs of force and is at a -15"

OpenStudy (anonymous):

"27lbs of force at 18degrees"

OpenStudy (anonymous):

Change to x,y component form.

OpenStudy (anonymous):

soo, <27,18> and <24,-15>

OpenStudy (anonymous):

are the vectors?

OpenStudy (anonymous):

No, those are not x, y components. They are magnitude and angle, or the polar r, theta components.

OpenStudy (anonymous):

\[ \langle x,y\rangle =\langle r\cos\theta ,r\sin\theta\rangle \]

OpenStudy (anonymous):

ok look at this for me, ill give an example

OpenStudy (anonymous):

component form: v_x = 18 cos (200) = -16.91 v_y = 18 sin (200) = -6.15 so, (-16.91,-6.15) yes?

OpenStudy (anonymous):

how did you get 200?

OpenStudy (anonymous):

it was just an example from another problem that i did... if -15 is a functional directional angle then i can find the component form

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so 24 cos(-15) 24 sin (-15) for x and y component form for john right?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

k thanks wio!!!

OpenStudy (anonymous):

so the answer should be <25.67, 8.34> ?

OpenStudy (anonymous):

Well, they may want you to convert it back to magnitude, angle.

OpenStudy (anonymous):

lol, how would you do that?